OpenStudy (anonymous):

Solve the quadratic equation y^2-7x-18=0

OpenStudy (anonymous):

\[y^2=7x+18\]\[y=\pm\sqrt{7x+18}\]

OpenStudy (anonymous):

\[\sqrt{7x+18}\]

OpenStudy (anonymous):

y=+-sqrt(7x+18) and x>-7/18

OpenStudy (solomonzelman):

you meant \[x^2-7x-18=0~~~~~~~~~~~~~~RIGHT?\]

OpenStudy (solomonzelman):

\[(x+2)(x-9)=0\]from there you can do it on your own.